OpenStudy (anonymous):

Three siblings collect rare coins. To determine the number of rare coins that Samantha has, take the number of rare coins kevin has, add 4, and then divide that sum by 2. To determine the number of rare coins Ben has, double the amount of rare coins Kevin has, subtract 4, and then multiply that difference by 2. How many rare coins does each sibling have if they have a total of 49 coins?

OpenStudy (ahsome):

Let use create 3 variables, one for each person and the number of rare coins: \(k=k\) \(s=\frac{k+4}{2}\) \(b=2(2k-4)\) Then add these these to one equation to find \(k\). \(k+b+s=49\) \(k+\frac{k+4}{2}+2(2k-4)=49\) Solve for \(k\), subsitute to the other equations to get \(b\) and \(s\)

OpenStudy (ahsome):

@paige11, do you understand?

OpenStudy (anonymous):

@Ahsome yes! but I don't understand the last part which is the last equation.. could u go over that part?

OpenStudy (ahsome):

I got to go, but someone else can help you. Help @iambatman, @arab